# CliBench: Multifaceted Evaluation of Large Language Models in Clinical Decisions on Diagnoses, Procedures, Lab Tests Orders and Prescriptions

### v1

@@ Appendix B Details of CliBench and experimental setup
B.1 Prompt examples
@@
B.2 Task target candidate ontology statistics
@@ XXII U00–U99 Codes for special purposes
B.4 Context lengths of LLMs
Model Maximum context length Radio of kept input after truncation (%)
Mean Medium Minimum
Mistral series 32768 98.58 100.00 24.84
LLaMA2 series 4096 53.82 48.90 2.50
LLaMA3 series 8192 87.20 100.00 6.68
B.5 Definition of first batch decisions
Among all the procedures, lab test orders, and prescriptions included in the structured database, we apply a filter and keep only the first batch of decisions as our target prediction ground truth. We define the "first batch" as the decision made at the first timestamp, appearing within 24 hours after the admission time for each kind of decision (procedures, lab test orders, prescriptions). Full lab test results are used as input for the diagnosis task, while only the first batch decisions are used as ground truth for the tasks except for diagnosis decisions.
B.6 Truncation rules
Appendix C Additional experimental results and analysis
C.1 Performance vs task difficulties
Performance vs patient length of stay. We demonstrate breakdown performance for admissions with various durations in Figure 4. We observe the both GPT-4o and LLaMA3 70B Instruct models produce worse recall when the duration increases, as the longer the stay, the more complicated the diagnosis decisions to make.
Figure 4: The diagnosis decision performance for full code level of GPT-4o and LLaMA3 70B Instruct by group of patients’ length of stay in terms of days.
Performance vs diversity of diagnoses. We present the breakdown performance of admissions subsets with different numbers of unique diagnosis chapters in Figure 5. The number of unique chapters represents the diversity and scope of diagnosis involved in the ground-truth billing code. The precision increases while a wider range of unique chapters is involved.
Figure 5: The diagnosis decision performance for full code level of GPT-4o and LLaMA3 70B Instruct by the number of unique diagnosis chapters in ground-truth billing diagnosis codes.
Performance vs number of ground-truth diagnoses. In Figure 6, we show the performance trend for admission instances with different numbers of ground-truth diagnoses. With the increasing of diagnosis numbers, the recall is getting lower while the precision is getting higher. From qualitative examples, we observe the models generate similar amounts of predictions no matter the complexity of the patient case. With more ground-truth diagnoses, it is easier to have a match, leading to improved precision.
Figure 6: The diagnosis decision performance for full code level of GPT-4o and LLaMA3 70B Instruct by number of diagnosis codes in ground-truth billing diagnosis codes.
C.2 Full results for diagnosis decisions
@@ C.5 Full results for prescriptions
C.6 Qualitative analysis and output examples
Flan-T5 XL:
Lung adenocarcinoma
Mistral v0.3:
Mistral Instruct v0.3:
BioMistral DARE:
LLaMA3:
LLaMA3 Instruct 8B:
LLaMA3 Instruct 8B SFT:
OpenBioLLM 8B:
LLaMA3 Instruct 70B:
GPT4o:
@@ Appendix D Limitations and potential negative impact
We would like to raise awareness that there might be miscoded diagnosis codes in the patient records. The billing ICD diagnosis codes are used as “ground-truth” diagnosis decisions to train our model and evaluate the performance for diagnosis prediction. The billing diagnosis codes do not exactly match the clinician’s diagnosis decisions, and they are input after the diagnoses are made. We acknowledge that the diagnosis code extracted from the EHR dataset should not be considered the best/perfect diagnosis decision. We also raise the potential data distribution issue as the training and evaluation data used in this work is largely collected for patients with ICU stay history. Thus, the evaluation result does not represent the generalized diagnosis prediction capability, and the trained model may yield compromised performance when different kinds of patient records are queried.

### v2

@@ Appendix B Details of CliBench and experimental setup
B.1 Summarized task definitions
We use Table 6 to summarize task definitions for all four clinical decisions.
Table 6: Clinical decisions task definition summary and comparisons.
Discharge diagnoses Procedures Lab test orderings Prescriptions
Description Diagnoses in the entire span of the patient admission given all patient records within this admission Initial (first batch) procedures to implement after the patient is admitted Initial (first batch) lab items to be conducted after the patient is admitted Initial (first batch) medications to be prescribed for the patient after the patient is admitted
Input Patient profile, medical record at admission, lab test results within the admission, radiology results within the admission, history diagnoses Patient profile, medical record at admission Patient profile, medical record at admission Patient profile, medical record at admission
Ground-truth A set of International Classification of Diseases, tenth Revision, Clinical Modification (ICD-10-CM) codes/concepts A set of ICD-10-Procedure Coding System codes/concepts A set of Logical Observation Identifiers Names and Codes (LOINC) codes/concepts A set of Anatomical Therapeutic Chemical (ATC) codes/concepts
Example (one item of the set) E10.618 (Type 1 diabetes mellitus with other diabetic arthropathy) B233Y0Z (Magnetic Resonance Imaging (MRI) of Multiple Coronary Artery Bypass Grafts using Other Contrast, Unenhanced and Enhanced) LP399135-5 (Direct antiglobulin test.IgA specific reagent | Red Blood Cells | Blood bank) C08DA01 (Verapamil)
B.2 Prompt examples
@@
B.3 Task target candidate ontology statistics
@@ B.4 ICD-10-CM chapters
B.5 Context lengths of LLMs
Model Maximum context length Radio of kept input after truncation (%)
Mean Medium Minimum
Mistral series 32768 98.58 100.00 24.84
LLaMA2 series 4096 53.82 48.90 2.50
LLaMA3 series 8192 87.20 100.00 6.68
B.6 Definition of first batch decisions
Among all the procedures, lab test orders, and prescriptions included in the structured database, we apply a filter and keep only the first batch of decisions as our target prediction ground truth. We define the “first batch” as the decision made at the first timestamp, appearing within 24 hours after the admission time for each kind of decision (procedures, lab test orders, prescriptions). Full lab test results are used as input for the diagnosis task, while only the first batch decisions are used as ground truth for the tasks except for diagnosis decisions.
B.7 Truncation rules
Appendix C Additional experimental results and analysis
C.1 Full results for diagnosis decisions
C.2 Full results for procedure decisions
C.3 Full results for lab test orders
C.4 Full results for prescriptions
C.5 Qualitative analysis and output examples
Flan-T5 XL:
Lung adenocarcinoma
Mistral v0.3:
Mistral Instruct v0.3:
BioMistral DARE:
LLaMA3:
LLaMA3 Instruct 8B:
LLaMA3 Instruct 8B SFT:
OpenBioLLM 8B:
LLaMA3 Instruct 70B:
GPT4o:
@@ XXII U00–U99 Codes for special purposes
14 LLaMA3.1 Instruct 8B 59.09 68.99 63.66 35.04 44.48 39.20 23.68 30.03 26.48
15 LLaMA3.2 Instruct 1B 58.03 52.61 55.19 30.19 27.87 28.98 17.58 16.12 16.82
16 LLaMA3.2 Instruct 3B 59.40 63.24 61.26 31.72 36.42 33.90 19.43 22.48 20.85
@@
14 LLaMA3.1 Instruct 8B 12.14 15.75 13.71 8.85 11.42 9.97 27.76 34.14 30.61
15 LLaMA3.2 Instruct 1B 8.26 7.66 7.95 5.54 5.15 5.34 23.92 21.88 22.86
16 LLaMA3.2 Instruct 3B 9.35 11.08 10.14 6.26 7.42 6.79 25.23 28.13 26.59
@@
14 LLaMA3.1 Instruct 8B 18.72 64.09 28.98 8.29 31.82 13.15 3.30 12.64 5.24
15 LLaMA3.2 Instruct 1B 17.35 50.21 25.79 7.71 23.17 11.57 2.85 8.43 4.26
16 LLaMA3.2 Instruct 3B 18.21 60.48 27.99 9.31 31.53 14.37 3.32 11.32 5.14
@@
14 LLaMA3.1 Instruct 8B 0.22 1.61 0.39 7.63 27.54 11.94
15 LLaMA3.2 Instruct 1B 0.15 0.77 0.26 7.02 20.64 10.47
16 LLaMA3.2 Instruct 3B 0.14 0.88 0.24 7.74 26.06 11.93
@@
14 LLaMA3.1 Instruct 8B 100.00 99.81 99.91 74.10 81.01 77.40 47.82 49.56 48.67
15 LLaMA3.2 Instruct 1B 99.91 99.81 99.86 75.63 77.25 76.43 49.57 40.89 44.82
16 LLaMA3.2 Instruct 3B 100.00 99.81 99.91 73.62 81.14 77.20 46.80 50.08 48.38
@@
14 LLaMA3.1 Instruct 8B 28.26 15.31 19.86 62.55 61.42 61.46
15 LLaMA3.2 Instruct 1B 31.34 10.65 15.90 64.11 57.15 59.25
16 LLaMA3.2 Instruct 3B 32.33 16.48 21.83 63.19 61.88 61.83
@@
14 LLaMA3.1 Instruct 8B 79.37 79.26 79.32 77.19 76.30 76.74 48.97 52.13 50.50
15 LLaMA3.2 Instruct 1B 79.16 68.31 73.34 77.30 63.59 69.78 48.16 39.56 43.44
16 LLaMA3.2 Instruct 3B 77.24 74.89 76.05 75.12 70.94 72.97 44.57 45.65 45.10
@@
14 LLaMA3.1 Instruct 8B 39.63 40.90 40.26 61.29 62.15 61.70
15 LLaMA3.2 Instruct 1B 38.50 30.43 33.99 60.78 50.47 55.14
16 LLaMA3.2 Instruct 3B 34.18 34.21 34.20 57.78 56.42 57.08
@@ Appendix D Limitations and potential negative impact
We would like to raise awareness that there might be miscoded diagnosis codes in the patient records. The billing ICD diagnosis codes are used as “ground-truth” diagnosis decisions to train our model and evaluate the performance for diagnosis prediction. The billing diagnosis codes do not exactly match the clinician’s diagnosis decisions, and they are input after the diagnoses are made. We acknowledge that the diagnosis code extracted from the EHR dataset should not be considered the best/perfect diagnosis decision. We also raise the potential data distribution issue as the training and evaluation data used in this work is largely collected for patients with ICU stay history. Thus, the evaluation result does not represent the generalized diagnosis prediction capability, and the trained model may yield compromised performance when different kinds of patient records are queried. There are potential mapping errors from the model response to a candidate code using the sentence BERT model, which could introduce noise in the evaluation results. Those seeking to use this benchmark should be cautioned that it is not suitable for picking up subtle changes in performance as evidence of a model’s superiority.